OpenStudy (anonymous):

Choose the equation of the line passing through the point (6, -2) and parallel to y = x + 7.

OpenStudy (anonymous):

y = x + 8 y = x - 4 y = x + 4 y = x - 8

OpenStudy (souvik):

the line parallel to y=x+7 has an equation y=x+k where k is some constant... now put x=6 and y=-2 and find k..

OpenStudy (anonymous):

slope of y=x+7 is 1 so the line passing throw (6,-2) is parallel to y=x+7 so ,slopes are equal m1=m2 now equation of a line passing through (6,-2) and slope 1 is (y-y1)=m(x-x1) y+2=1(x-6) x-y-8=0

OpenStudy (anonymous):

I don't under stand

OpenStudy (anonymous):

so answer is y=x-8

OpenStudy (anonymous):

yeah but I don't get how you did it

OpenStudy (souvik):

parallel lines have same slope....right?@lexi123479

OpenStudy (anonymous):

as the line y=x+7 and the line we have to find are parallel they have same slopes

OpenStudy (anonymous):

so find the slope of y=x+7

OpenStudy (anonymous):

hence it is the slope of the line we have to find

OpenStudy (anonymous):

so we got the slope and point is given with point slope form find the equation

OpenStudy (anonymous):

did you get it